# LEAM: A Prompt-only Large Language Model-enabled Antenna Modeling Method

### v1

@@ Abstract
Antenna modeling is a time-consuming and complex process, decreasing the speed of antenna analysis and design. In this paper, a large language model (LLM)- enabled antenna modeling method, called LEAM, is presented to address this challenge. LEAM enables automatic antenna model generation based on language descriptions via prompt input, images, descriptions from academic papers, patents, and technical reports (either one or multiple). The effectiveness of LEAM is demonstrated by three examples: a Vivaldi antenna generated from a complete user description, a slotted patch antenna generated from an incomplete user description and the operating frequency, and a monopole slotted antenna generated from images and descriptions scanned from the literature. For all the examples, correct antenna models are generated in a few minutes. The code can be accessed via https://github.com/TaoWu974/LEAM.
Index Terms:
Large language model, antenna, electronic design automation, GPT, antenna modeling, human–computer interaction
@@ I Introduction
In antenna research and applications, there is a demanding need for exploring novel antenna structures and reproducing designs from the literature. Antenna modeling is often unavoidable in this process, costing a lot of time and effort. Depending on the user’s experience, modeling a complex antenna could take from several tens of minutes to hours, and the number of such antenna modeling tasks grows rapidly with the growth of the antenna industry. It would be ideal to reduce or remove this repetitive and error-prone work. This inspires us to automate antenna modeling, speeding up the antenna design and innovation process.
@@
It is routine to use the graphical user interface (GUI) in electromagnetic (EM) modeling and simulation tools for antenna modeling and visualization. These visualized antenna structures can be represented in macro files, which are command lines describing the model. For example, CST Microwave Studio employs Visual Basic for Applications (VBA) macros, and HFSS supports IronPython macros. Hence, by generating appropriate macro files, the antenna modeling process can become automated.
However, macro files are complex, and manual scripting is even more time-consuming than drawing the model using GUI. For example, importing material to an object is a straightforward click using the GUI of EM simulation tools, but tens of parameters are needed in the macro file. The gap between the simple language description of the user and the complex macro file that can correctly define the antenna model becomes the bottleneck, which has not been solved till now. However, a recent innovation, Large Language Models (LLMs), provides a promising solution, bridging simple descriptions in prompt (a prompt is the text input to the LLM) and complex macro files by its code generation [1], geometry understanding [2], parameterization [3] and learning ability.
@@
Although have not been used in antenna modeling, LLMs are investigated for automated macro scripting for 3D mechanical structure modeling. Some pioneer research works and findings are as follows. [4] shows weak prompts can lead to poor responses and task failures. [5] shows that LLMs with vision capabilities (i.e., image recognition) outperform text-only LLMs as the modeling complexity increases. [6] collects macro templates and benchmark test cases to assess LLM-based modeling techniques. SolidWorks employs LLMs in modeling mechanical components [7].
@@
Despite the advancements in applying LLMs to 3D modeling, two limitations are restricting their applications in antenna modeling: (1) There is limited research considering multi-modal input (i.e., not only textual descriptions, but also image input) [5], but is highly needed in LLM-enabled antenna modeling, especially for reproducing models from the literature. (2) Antenna modeling often consists of a series of solids’ modeling and boolean operations among them, which is quite different from mechanical modeling’s process of selecting one model and configuring its parameters [5].
@@
To address these challenges, the LLM-enabled antenna modeling method (LEAM) is proposed. The macro scripting process is divided into systematic steps, and a set of LLM tools based on OpenAI LLM Application Programming Interface (APIs) with designated prompts including responses from other LLM tools are designed for these steps. The input of LEAM is textual descriptions and images (if available), and the output is the antenna model. Currently, CST Microwave Studio is used as the EM simulator. To the best of our knowledge, this research is the first work investigating LLM-enabled antenna automated modeling. Moreover, LEAM is a prompt-only method, avoiding complex and time-consuming fine-tuning (i.e., additional training using specific data for antennas) or retrieval-augmented generation (i.e., providing knowledge databases, such as documents, and books), improving usability.
II The LEAM Methodology
II-A The workflow of LEAM
@@
In LEAM, the antenna modeling approach is divided into systematic steps, and for each step, a tool is created. The tools are shown in Fig. 1 and Table I. The user employs the tools mostly sequentially, except in the first step, there is selection, which will be detailed later.
@@
UpdatePara *.bas UpdatePara.bas o1
RunVBA *.bas Model.cst N/A
As shown in Fig. 1, the modeling process is divided into four stages, which are identifying solids, defining solids, modeling solids, and executing macros. (In this work, CST Microwave Studio is used, where solid is the fundamental modeling unit.) In the first stage, if the provided description is sufficient, the strong descriptions to solids (Strong D2S) tool will be used; otherwise, the weak descriptions to solids (Weak D2S) tool will be used. Both tools have the same output, which lists all solids with their materials, types (e.g., substrate, feeding line, slot, etc.), geometric shapes (e.g., brick, cylinder, extruded 2D shape, etc.), and positions.
@@
The third stage is to model solids, which can be divided into 3D shapes or 2D+ shapes. Given the solid list with dimensions, parameters, and materials, the Model3D tool identifies and models the 3D shapes, and produces the 3D_Model.bas file. 2D+ shape refers to extruded or rotated 2D shapes. The Model2D+ tool first models closed 2D shapes and then extrudes and/or rotates these 2D shapes into 3D shapes, and produces the 2D+_Model.bas file. Then, given the modeled solids in the above bas files and the solid list with dimensions, the Boolean tool determines sequential boolean operations. For example, adding the connected solids, subtracting the slot solids.
The last stage is executing macro files. The RunVBA tool executes all bas files obtained before, including Para.bas, Materials.bas, 3D_Model.bas, 2D+_Model.bas, and Boolean.bas. These macros are saved to the history list of CST Microwave Studio. The automated antenna modeling is completed. After this step, an optional tool is the UpdatePara tool, which is used to change parameters. When the user provides new parameter sets, the UpdatePara.bas is generated and executed to update the antenna model.
@@ II-B LLM tool creation
As said above, 9 LLM tools and a RunVBA tool are defined in LEAM. Fig. LABEL:fig:_tool_I/O shows a tool structure using the Dimension tool as an example. Each tool accepts pre-defined input, input from former tools, and user input. The pre-defined input is the major effort to link LLM with antenna modeling, including role and task prompts, functional instructions, and knowledge files. The role and task prompts provide a task-oriented identity to help the LLM align with task requirements [8]. The functional instructions define the task details and output rules (e.g., ”Output VBA code only”) [9]. Knowledge files are examples of macros and formatted output to guide the LLM.
Figure 2: The Dimension tool
It is worth noting that various LLM models are carefully selected for each tool. With the advancement of reasoning LLMs such as OpenAI o1 [10] and Deepseek R1 [11], LLMs can now build more complex workflows using the Chain of Thought (CoT) [10], aligning seamlessly with complex scripting requirements of antenna modeling. Our pilot experiments show that OpenAI o1 is better at understanding a topology and dismantling it into basic geometric shapes, while GPT-4o outperforms OpenAI o1 in image recognition. Hence, to trade off the reasoning ability and image recognition ability, the general principle is that the base model (GPT-4o-2024-11-20) is used for most multi-modal input and the reasoning model (o1-2024-12-17) is used for text-only input. However, when advanced reasoning is critical, such as the D2S tool, OpenAI o1 is employed even though the input is multi-modal.
@@ II-C Demonstration of LEAM by a simple case
Figure 3: Example: L-slot on the patch. $Patch_{L}$ = 8, $Patch_{W}$ = 10, $Slot1_{L}$ = 4, $Slot1_{W}$ = 1, $Slot2_{L}$ = 6, $Slot2_{W}$ = 1, $Slot1_{offset}$ = 2. (unit: mm.)
@@
Start from [$\mathrm{Slot_{Offset}}$,$\mathrm{Slot_{Offset}}$]
To [$\mathrm{Slot_{Offset}+Slot2_{L}}$, $\mathrm{Slot_{Offset}}$]
To [$\mathrm{Slot_{Offset}+Slot2_{L}}$, $\mathrm{Slot_{Offset}+Slot2_{W}}$]
To [$\mathrm{Slot_{Offset}+Slot1_{W}}$, $\mathrm{Slot_{Offset}+Slot2_{W}}$]
To [$\mathrm{Slot_{Offset}+Slot1_{W}}$, $\mathrm{Slot_{Offset}+Slot1_{L}}$]
To [$\mathrm{Slot_{Offset}}$, $\mathrm{Slot_{Offset}+Slot1_{L}}$]
To [$\mathrm{Slot_{Offset}}$,$\mathrm{Slot_{Offset}}$]
@@ III Experimental Results
In this section, three case studies are used to demonstrate LEAM. The first case aims to model a Vivaldi antenna with detailed description text input by the user. The second case aims to model a patch antenna operating at 2.45 GHz but the description text input by the user is incomplete. The third case aims to reproduce the model of a monopole antenna using descriptions and images scanned from [12]. For all the test cases, LEAM successfully generated the correct antenna model in a few minutes.
III-A Case 1: The Vivaldi antenna modeling
[13] indicates that the exponential non-uniform slot plays an important role in reducing the antenna’s physical dimension, and the circular stub on the back of the substrate provides better impedance matching for a smooth microstrip-to-slotline transition. Inspired by this, a Vivaldi antenna is proposed with a substrate (30 mm $\times$ 20 mm), a slotline defined by two symmetric splines and a circular slot on the patch above the substrate, and a feeding structure of two connected feedlines with a circular stub under the substrate. The splines are defined by 20 points, whose coordinates are as follows.
$Left$ $:(x_{i},y_{i}),\quad\forall i=1,2,\dots,20$ (1)
$Right$ $:(30-x_{i},y_{i})\quad\forall i=1,2,\dots,20$
@@ III-A Case 1: The Vivaldi antenna modeling
A complete text description of the Vivaldi antenna is prepared. The description for the slot curve of the patch is as follows.
  • Model two symmetric splines and connect their top and bottom points to form a closed shape to extrude.
  • The left-spline is defined by 20 points, whose y axis are 20, 19, … 1, and the X axis are variables. (Here we will have 20 variables).
  • The point’s $x$ axis is ascending within the range (0, 15), exclusive. Meanwhile, the right half’s $x$ value is 30-$x\_i$.
Using the complete text description for the Strong D2S Tool, 30 parameters and 2 materials are imported. 6 3D solids (substrate, front patch, front cylinder, back rectangular#1, back rectangular#2, back cylinder) and a 2D+ solid (slotline) are modeled and shown in Fig. 4, whose parameters are updated by the UpdatePara tool to form the Vivaldi shape. The generated antenna model is shown in Fig. 4.
(a) Front view
(b) Back view
Figure 4: Vivaldi antenna. SL = 30, SW = 20, $r_{1}=2$, $gap_{1}=1$, $L_{1}=7$, $L_{2}=8$, $gap_{2}=5$, $W_{1}=2$, $W_{2}=1$, $r_{1}=2$. (unit: mm.)
III-B Case 2: The patch antenna modeling
Unlike case 1, the description for case 2 is incomplete. The only prompt is ”I want to design a rectangular-patch antenna with a rectangular slot working at 2.45 GHz”. The Weak D2S tool in LEAM identifies five 3D solids (substrate, patch, slot, feedline, and ground plane) with 2 materials and 4 parameters. The modeled antenna is shown in Fig. 5, with its $S_{11}$ response exactly at 2.45 GHz shown in Fig. 6. To identify the appropriate dimensions of the patch, LEAM builds CoT and applies the transmission line model in [14]. This case demonstrates that even without a complete description, LEAM can make use of CoT in LLM technology and help users who are not antenna experts.
Figure 5: Slotted patch antenna @ 2.45GHz. PatchW = 38, PatchL = 28, SlotW = 2, SlotL = 10. (unit: mm.)
Figure 6: Simulated $S_{11}$ response for the slotted patch antenna.
III-C Case 3: The monopole antenna modeling
The third case is a slotted monopole antenna from the literature. Its description, topology figure, and parameter list figure are scanned from [12] (Fig. 7(a)). The strong D2S tool identifies 7 3D solids (substrate, circular patch, horizontal slot, vertical slot, left ground plane, and right ground plane) with 12 parameters (10 parameters in the figure and 2 thickness parameters in the text) and 2 materials. Moreover, we add a prompt to the Parameter tool to correct a typo in [12] ($S_{L}=M_{L}+DP_{R}+0.2$). The modeled antenna is shown in Fig. 7.
(a) Descriptions of the slotted monopole antenna, retrieved from Section IV, second paragraph of [12].
(b) Monopole
Figure 7: Slotted monopole antenna. $DP_{R}$ = 6.58, $S_{W}$ = 13.43, $SL_{T}$ = 1, $SL_{V}$ = 7.9, $SL_{H}$ = 7.9, $M_{L}$ = 25.08, $RP_{L}$ = 6.67, $M_{W}$ = 1.2, $M_{G}$ = 0.3, $S_{L}$ = 31.86, $RP_{W}$ = 5.815. (unit: mm.)

### v2

# Large Language Model-Based Intelligent Antenna Design System

@@ Abstract
Antenna simulation typically involves modeling and optimization, which are time-consuming and labor‑intensive, slowing down antenna analysis and design. This paper presents a prototype of a large language model (LLM)-based antenna design system (LADS) to assist in antenna simulation. LADS generates antenna models with textual descriptions and images extracted from academic papers, patents, and technical reports (either one or multiple), and it interacts with engineers to iteratively refine the designs. After that, LADS configures and runs an optimizer to meet the design specifications. The effectiveness of LADS is demonstrated by a monopole slotted antenna generated from images and descriptions from the literature. To improve gain stability across the 3.1–10.6 GHz ultra-wide band, LADS modifies the cross-slot into an H-slot and changes substrate material, followed by parameter optimization. As a result, the gain variation is reduced while maintaining the same gain level. The LLM-enabled antenna modeling (LEAM) is available at: https://github.com/TaoWu974/LEAM.
@@ I Introduction
In antenna research and applications, there is a demanding need for exploring novel antenna structures and reproducing designs from the literature. Antenna modeling and optimization are often unavoidable in this process, costing a lot of time and effort. Depending on the user’s experience, modeling and optimizing a complex antenna could take from dozens of hours to days, and the demand for such simulations continues to rise with the growth of the antenna industry. Therefore, it is desirable to reduce repetitive and error-prone modeling work and employ the human-free optimization. This motivates the automation of antenna modeling and optimization, speeding up the antenna design and innovation process.
@@
It is common practice to use the graphical user interface (GUI) in electromagnetic (EM) simulation tools for antenna modeling and visualization first, followed by implementing external optimization scripts, typically in Python or MATLAB, to control the simulation and tune the design parameters. The visualized antenna structures can be represented in macro files, which are command lines describing the model. Macros also act as an interface between external optimization scripts and EM simulation tools to enable automated control of the simulations. For example, CST Microwave Studio employs Visual Basic for Applications (VBA) macros, and HFSS supports IronPython macros. To enable automation, two types of macro files are typically required: one for constructing the antenna model and another for controlling the simulation flow and parameter updates during optimization.
However, modeling macro files are complex, and manual scripting is even more time-consuming than drawing the model using GUI. For example, importing material to an object is a straightforward click using the GUI of EM simulation tools, but tens of parameters are needed in the macro file. The gap between the simple language description of the user and the complex macro file that can correctly define the antenna model becomes the bottleneck, which has not been solved till now. Moreover, even with existing optimization scripts, it remains nontrivial for antenna engineers to configure optimization, including setting optimizer parameters and preparing macros to link simulation tools and optimization scripts. For example, design variables and their search ranges must be defined correctly considering geometry constraints, while from a view of optimization goals, constraints and objectives need to be handled separately. These challenges make it difficult for designers to obtain optimal parameters efficiently, and the misconfigurations can easily lead to optimization failures.
@@
A recent innovation, Large Language Models (LLMs), provides a promising solution, bridging prompt and codes from macros and optimization scripts by its ability of geometry understanding [parra2024can], parameterization [wu2023cad] and code generation [lin2024llm]. Prompt is the input to the LLMs and can be texts and/or images. Although have not been used in antenna design, LLMs are investigated for automated macro scripting for 3D mechanical structure modeling. Some pioneer research works and findings are as follows. [zamfirescu2023johnny] shows weak prompts can lead to poor responses and task failures. [li2025llm4cad] shows that LLMs with vision capabilities (i.e., image recognition) outperform text-only LLMs as the modeling complexity increases. [badagabettu2024query2cad] collects macro templates and benchmark test cases to assess LLM-based modeling techniques. SolidWorks employs LLMs in modeling mechanical components [deng2024investigation].
Despite the advancements in applying LLMs to 3D modeling, three limitations are restricting their applications in antenna modeling and optimization:
@@
  • There is limited research considering multi-modal input (i.e., not only textual descriptions, but also image input) [li2025llm4cad], but is crucial for LLM-enabled antenna modeling, especially for reproducing models from the literature.
  • Antenna modeling typically involves creating multiple solids and performing Boolean operations among them, which differs from the mechanical modeling process of selecting one model and configuring its parameters [li2025llm4cad].
  • To date, there has been little research aimed at integrating antenna modeling and optimization into a unified workflow.
@@
To address these challenges, the LLM-based antenna design system (LADS) is proposed, which consists of an LLM-enabled antenna modeling method (LEAM) and LLM-controlled optimization. The modeling macro scripting process is divided into systematic steps, and a set of LLM tools based on OpenAI LLM Application Programming Interface (APIs) with designated prompts including responses from other LLM tools are designed for these steps. The input of LEAM is textual descriptions and images (if available), and the output is the antenna model. Currently, CST Microwave Studio is used as the EM simulator. To the best of our knowledge, this is the first study to investigate LLM-enabled antenna automated modeling. After that, human engineers can propose new design goals to LLM, then review and select solutions from LLMs to refine the antenna design. Finally, LLM configures and runs parameter optimization using SB-SADEA, a self-adaptive Bayesian neural network surrogate-model-assisted differential evolution for antenna optimization [liu2022efficient].
II The System Overview and Methodology
II-A The LADS workflow
@@
In LADS, the antenna modeling and optimization are divided into systematic steps, and for each step, several tools are created. The tools are shown in Fig. 1 and Table I. Users typically employ the tools sequentially, except for the first step, there is selection, which will be detailed later.
@@
Modeling *.bas Model.cst N/A
ParaControl *.txt & Para.bas Simu.bas o1
Interface Simu.bas Interface.yml o1
ConfigOpt Text & Interface.yml Opt_config.yml o1
SB-SADEA *.yml Optimal.cst N/A
As shown in Fig. 1, the design process is divided into four stages, which are identifying solids, defining solids, executing macros, and optimization. (In this work, CST Microwave Studio (CST MWS) is used, where solid is the fundamental modeling unit.) In the first stage, if the provided description is sufficient, the strong descriptions to solids (Strong D2S) tool will be used; otherwise, the weak descriptions to solids (Weak D2S) tool will be used. Both tools have the same output, which lists all solids with their materials, types (e.g., substrate, feeding line, slot, etc.), geometric shapes (e.g., brick, cylinder, extruded 2D shape, etc.), and positions. The Strong D2S tends to translate the user’s instructions only, while Weak D2S autofills the incomplete descriptions with LLM’s knowledge about antennas.
@@
After that, 3D and 2D+ solids modeling macros are generated separately, followed by boolean operations. Given the solid list with dimensions, parameters, and materials, the Model3D tool identifies and models the 3D shapes, and produces the 3D_Model.bas file. 2D+ shape refers to extruded or rotated 2D shapes. The Model2D+ tool first models closed 2D shapes and then extrudes and/or rotates these 2D shapes into 3D shapes, and produces the 2D+_Model.bas file. Then, given the modeled solids in the above bas files and the solid list with dimensions, the Boolean tool determines sequential boolean operations in the Boolean.bas file. For example, adding the connected solids, subtracting the slot solids.
The third stage is to execute macro files. The Modeling tool executes all bas files obtained previously, including Para.bas, Materials.bas, 3D_Model.bas, 2D+_Model.bas, and Boolean.bas. These macros are saved to the history list of CST Microwave Studio. The automated antenna modeling is completed. After that, ParaControl tool determines the parameters to be optimized and generates the corresponding parameter control and simulation bas file Simu.bas.
Figure 2: The LLM-based Dimension tool I/O
Finally, the Interface tool reads in the Simu.bas file to generate an interface.yml file to configure the interface of controlling the design variables and calling CST MWS to run EM simulations, while the ConfigOpt tool creates an Opt_config.yml file to configure the optimization, including bounds, constrains, objectives and other optimization settings. Finally, SB-SADEA [liu2022efficient] runs the optimization of antenna parameters with the given *.yml files, and outputs an optimal design.
@@ II-B LLM tool creation
As said above, bas executing tool, SB-SADEA optimizer tool and 11 LLM tools are defined in LADS. Fig. 2 shows a tool structure using the Dimension tool as an example. Each tool accepts pre-defined input, input from former tools, and user input. The pre-defined input is the major effort to link LLM with antenna modeling, including role and task prompts, functional instructions, and knowledge files. The role and task prompts provide a task-oriented identity to help the LLM align with task requirements [zheng2023helpful]. The functional instructions define the task details and output rules (e.g., ”Output VBA code only”) [zhang2024instruct]. Knowledge files are examples of macros and formatted output to guide the LLM.
It is worth noting that various LLM models are carefully selected for each tool. With the advancement of reasoning LLMs such as OpenAI o1 [jaech2024openai] and Deepseek R1 [guo2025deepseek], LLMs can now build more complex workflows using the Chain of Thought (CoT) [jaech2024openai], aligning seamlessly with complex scripting requirements of antenna modeling. Our pilot experiments show that OpenAI o1 is better at understanding a topology and dismantling it into basic geometric shapes, while GPT-4o outperforms OpenAI o1 in image recognition. Hence, to trade off the reasoning ability and image recognition ability, the general principle is that the base model (GPT-4o-2024-11-20) is used for most multi-modal input and the reasoning model (o1-2024-12-17) is used for text-only input. However, when advanced reasoning is critical, such as the D2S tool, OpenAI o1 is employed even though the input is multi-modal.
@@ II-C Demonstration of LEAM by a simple case
Figure 3: Example: L-slot on the patch. $Patch_{L}$ = 8, $Patch_{W}$ = 10, $Slot1_{L}$ = 4, $Slot1_{W}$ = 1, $Slot2_{L}$ = 6, $Slot2_{W}$ = 1, $Slot1_{offset}$ = 2. (unit: $m$.)
@@
Start from $\mathrm{Slot_{Offset}}$$\mathrm{Slot_{Offset}}$]
To $\mathrm{Slot_{Offset}+Slot2_{L}}$, $\mathrm{Slot_{Offset}}$]
To $\mathrm{Slot_{Offset}+Slot2_{L}}$, $\mathrm{Slot_{Offset}+Slot2_{W}}$]
To $\mathrm{Slot_{Offset}+Slot1_{W}}$, $\mathrm{Slot_{Offset}+Slot2_{W}}$]
To $\mathrm{Slot_{Offset}+Slot1_{W}}$, $\mathrm{Slot_{Offset}+Slot1_{L}}$]
To $\mathrm{Slot_{Offset}}$, $\mathrm{Slot_{Offset}+Slot1_{L}}$]
To $\mathrm{Slot_{Offset}}$$\mathrm{Slot_{Offset}}$]
@@ III Experimental Results
In this section, an ultra-wide band (UWB) slotted monopole antenna’s modeling and optimization are used to demonstrate LADS, and some other cases of modeling method LEAM can accessed by https://github.com/TaoWu974/LEAM. The slotted patch antenna model is reproduced using descriptions and images scanned from [liu2022efficient].
The description, topology figure, and parameter list figure are scanned from [liu2022efficient] (Fig. 4). The strong D2S tool identifies seven 3D solids (substrate, circular patch, horizontal slot, vertical slot, left ground plane, and right ground plane) with 12 parameters (10 parameters from the paper’s parameter list and 2 thickness parameters in the text) and 2 materials, FR-4 for the substrate and copper for the patch. A typo in [liu2022efficient] is corrected by adding a prompt to the Parameter tool ($S_{L}=M_{L}+DP_{R}+0.2$). Then, to refine the gain stability over the frequency band from 3.1 GHz to 10.6 GHz, we pass this requirement, identified solids, and descriptions from the paper as a prompt to the LLM. Ten solutions are given, and two of them are chosen by the engineer: 1) Change the substrate material from FR-4 to Rogers RT 5880, 2) Change the slot structure from the cross-slot to the H-slot. Some other unselected solutions include adding parasitic element, increasing substrate thickness, and applying a taper feeding.
(a) Descriptions
(b) Topology
Figure 4: Cross-slotted monopole antenna, retrieved from Section IV of [liu2022efficient].
TABLE II: Design Variables of the H-Slotted Monopole Antenna
Design Variable Design Variable
Circle patch radius ($\mathrm{DP}_{R}$) Substrate width ($\mathrm{S}_{W}$)
Horizontal slot width ($\mathrm{SW}_{H}$) Vertical slot width ($\mathrm{SW}_{V}$)
Horizontal slot length ($\mathrm{SL}_{H}$) Vertical slot length ($\mathrm{SL}_{V}$)
Vertical slot offset ($\mathrm{S}_{O}$) Ground plane length ($\mathrm{G}_{L}$)
Microstrip width ($\mathrm{M}_{W}$) Microstrip gap ($\mathrm{M}_{G}$)
Feed guide width ($\mathrm{F}_{W}$) Microstrip length ($\mathrm{M}_{L}$)
Substrate length ($\mathrm{S}_{L}$) = $\mathrm{M}_{L}+\mathrm{DP}_{R}+0.2$ ($m$)
Ground plane width ($\mathrm{G}_{W}$) = $(\mathrm{S}_{W}-2\times\mathrm{M}_{G}-\mathrm{M}_{W})/2$ ($m$)
As a result, the modeled antenna is shown in Fig. 5, and 12 design variables are summarized in Table II. The antenna’s substrate is of Rogers RT 5880 with a thickness of 0.8 mm. It is noteworthy that the solution space of the H-slot topology encompasses that of the cross-slot structure. Specifically, when the horizontal offset between the two vertical slots is set to zero, or when their widths exceed the offset, the H-slot degenerates into a cross-slot configuration. Subsequently, SB-SADEA is employed to optimize these variables. The antenna is discretized with a mesh density of 15 cells per wavelength, resulting in approximately 75,000 mesh cells in total. Each EM simulation takes about 50 seconds to complete.
Figure 5: H-slotted monopole antenna
The optimization aims to minimize the variance of the realized gain ($G$) over the frequency range of 3.1–10.6 GHz, subject to two constraints: (1) the maximum reflection coefficient must remain below 10-10 dB, and (2) the minimum realized gain must exceed 1 dB. The optimization problem is formulated as follows:
$\min$ $\mathrm{Var}(G)$ (1)
@@
[…]
(a) Reflection coefficient
(b) Realized gain
Figure 6: The simulated results for the H-slotted antenna
TABLE III: Comparison of Proposed and Existing Antennas
Ref. $\epsilon_{r}$ Freq. (GHz) Gain (dBi) Volume $\lambda_{0}\times\lambda_{0}\times\lambda_{0}$
[…]
* This work
A design satisfying the two constraints is obtained after 256 EM simulations, followed by continuing optimization and obtaining the optimal design with 721 EM simulations using about 12 hours. Fig. 6 shows the antenna’s reflection coefficient and realized gain. Also, the obtained antenna is compared to existing UWB antennas in the literature in Table III, where $\lambda_{0}$ is the free-space wavelength at 3.1 GHz (96.8 mm). The realized gain range is reduced to $0.99\textit{ dBi}$ from $1.71\textit{ dBi}$ of the reference [liu2022efficient], showing improved gain stability.